Question 6

6. Solve the following inequality using the table of signs method. (r – 1)(x + 1)(r + 2) -2

can you explain? ty kindly
Answer #1


som find fw-(2-1) (a+1)(x+2) co (-2) - find the point where the ear can be zero the If x-1=0 x+1=0 2+2=0 . x-2=0 x1 m2-1 , x=